OpenStudy (anonymous):

Help/ check y = 2x - 3 would be y= x-3/2

OpenStudy (shamil98):

check what?... y = 2x -3 is already in slope-intercept form what are you asking for? .-.

OpenStudy (anonymous):

if my answer was right.

OpenStudy (mathlegend):

@marieeeloveee are you checking to see if it is the inverse?

OpenStudy (anonymous):

yes

OpenStudy (mathlegend):

At a glance I will have to say no.

OpenStudy (mathlegend):

In the original the 2 is being multiplied so you divided... which is wonderful! However, your 3 is being subtracted so in the inverse shouldn't it now have a plus in front?

OpenStudy (shamil98):

switch x and y and solve for y. x = 2y - 3 x + 3 = 2y x+3/2 = y

OpenStudy (shamil98):

(x+3)/2 = y i should say.

OpenStudy (mathlegend):

\[y ^{-1}=\frac{ x+3 }{ 2 }\]

OpenStudy (anonymous):

okay I see now, thank you @shamil98 and @MathLegend